OpenStudy (anonymous):

if 2(a/k)+1=9 , what does a/k equal?

OpenStudy (mathstudent55):

Treat a/k as a variable, and solve for a/k.

OpenStudy (anonymous):

2(a/k)+1=9 2(a/k)=8 (a/k)=8/2 (a/k)=4

OpenStudy (mathstudent55):

You may substitute y for a/k to make it look simpler. Let y = a/k, then 2y + 1 = 9 Solve for y.

OpenStudy (anonymous):

4 do it backwards, 9-1=8 8/2=4 a/k=4